A researcher orders a broth of 0.3% glucose for her lab. However she needs a stronger broth,one that is 26.4% glucose. Fortunately, she has 69.6 liters of 34.8% glucose broth in the stock room. How much 26.4% glucose broth can she make?


I need help with this problem, I believe it is solving for a percent mixture problem. Thanks.

Answer by ankor@dixie-net.com(12701)

A researcher orders a broth of 0.3% glucose; needs 26.4% glucose. has 69.6 liters of 34.8% glucose. How much 26.4% glucose broth can she make?
:
Let x = amt (in liters) of .3 % glucose required:
and
69.6 liters of 34.8% available
Therefore
Resulting amt = (69.6 + x) liter
:
.003x + .348(69.6) = .264(69.6+x)
:
.003x + 24.2208 = 18.3744 + .264x
:
.264x - .003x = 24.2208 - 18.3744
:
.261x = 5.8464
:
x = 22.4 liters of .03% glucose
Therefore
Total amt = 69.6 + 22.4 = 92 liters of 26.4% stuff
:
:
Check solution:
.003(22.4) + .348(69.6) = .264(92)
.0672 + 24.2208 = 24.288

e